OpenStudy (anonymous):

Which expression is the simplified form of the square root of m to the eighth power.? m4 m4the square root of m. m2 m2the square root of m.

OpenStudy (zehanz):

Whether it is \(\sqrt{m^8}\) or \((sqrt{m})^8\) doesn't matter, because the two are equal. Then maybe it is easier to deal with the second: \[(\sqrt{m})^8=\sqrt{m} \cdot \sqrt{m} \cdot \sqrt{m} \cdot \sqrt{m} \cdot \sqrt{m} \cdot \sqrt{m} \cdot \sqrt{m} \cdot \sqrt{m} =\]\[( \sqrt{m} \cdot \sqrt{m}) \cdot( \sqrt{m} \cdot \sqrt{m}) \cdot( \sqrt{m} \cdot \sqrt{m}) \cdot( \sqrt{m} \cdot \sqrt{m})=\]Do you know what is \[\sqrt{m} \cdot \sqrt{m}\]

OpenStudy (anonymous):

m^2?

OpenStudy (zehanz):

No, that is the answer of m*m

OpenStudy (zehanz):

What do you get when you do this:\[\sqrt{25}\cdot \sqrt{25}\]